# Chiral excitation flows of multinode network based on synthetic gauge fields

### v1

@@ VI.2 Emitters in magnetic photonic lattice
which manifests a long-range dipole-dipole interaction with complex hopping amplitudes $\tilde{J}_{jk}\sim g^{2}e^{i\theta_{jk}}/(\omega_{e}-\omega_{\ell=0})$ [52]. By setting the gauge invariant flux $\theta_{4}\equiv\theta_{12}+\theta_{23}+\theta_{34}+\theta_{41}=2\pi$ to break the TR symmetry of each triangular sub-structures, this effective dipole-dipole interaction can to some extent construct the four-node spin-aSGF model. Fig. 8(b) shows the chiral dynamics of excitation, where two factors contribute to the imperfection of this chiral flow. Firstly, in this configuration, the coupling parameter of the auxiliary emitter is $\beta_{c}=1.48$ instead of $2$, which is unable to generate a perfect four-node chiral flow. Secondly, the unavoidable NNN hopping between emitters from diagonal sites also reduces the fidelity of the chiral flow. However, by introducing external coupling channels [18, 21, 13, 70] to eliminate these effects, it is still possible to realize a perfect chiral flow between atoms in this system.
@@ VII Conclusion and Outlook
In conclusion, we investigate the boson-SGF model and show the chiral excitation flow in the three-node case with a broken TR symmetry. However, the pure SGF model cannot drive the chiral dynamics when it extends to the four-node case. By introducing the auxiliary emitter to couple to network nodes, we recreate the chiral flow in the four-node boson-aSGF network. The excitation dynamics of this four-node boson-aSGF network suggests that it can exhibit chiral flow only when the synthetic flux of each triangular sub-structure breaks TR symmetry, i.e., $\theta_{\vartriangle}\neq\mathbb{Z}\pi$. This links the well-studied three-node chiral flow and the boson-aSGF model for creating perfect chiral flows in multinode networks, where these triangular sub-structures are also retained, with $(\mathbb{Z}+1)\pi/2$ synthetic flux. When we generalize this framework to the spin or two-level atomic system where TR symmetry is preserved, it demonstrates opposite chiral dynamics for the spin-up and spin-down states. Further, we introduce a universal scheme for constructing the $n$-node chiral flow Hamiltonians with the aSGF model and provide illustrative examples for odd and even $n$, respectively. During this process, we find the criteria for a Hamiltonian to support the chiral flow in a specific excitation subspace, which indicates the requirement of a symmetric and equally spaced energy spectrum and a complete set of chiral modes. Our framework for finding chiral flows applies to various systems such as bosons, spins, and two-level atoms. It is also applicable to different types of interactions, including the NN complex hopping forming a two-body Hamiltonian, as well as the three-body interactions like SCI and ASI in spin systems. On the other hand, to realize the chiral flow in a larger network, the number of required hoppings scales as $n^{2}$, showing a significant complexity growth. Therefore, we propose a ladder network with the complexity scaling as $n$. Instead, it can serve as a more practical scheme for unidirectional transmission between remote emitters. In addition, the imperfections of networks are also discussed, where the fidelity of chiral flow shows robustness in the presence of hopping disorder but is sensitive to the inhomogeneity of node (boson mode or atomic) frequency.
@@
The multinode aSGF model and the extension of the four-node network can be realized with superconducting circuits by employing the Floquet modulation, which provides a new application to implement perfect state transfer in circuit QED systems. Moreover, the magnetic photonic lattice can also mimic this model to a certain extent with atomic emitters. The experimental feasibility, as well as the tolerance to imperfections, make our scheme practical and promising to realize chiral state transfer and unidirectional entanglement distribution. Also, we expect that our proposal could be utilized to develop non-reciprocal quantum devices [63] like $n$-node circulators and benefit future quantum technologies.
@@ VIII Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 12375025 and No. 11874432), the National Key R&D Program of China (Grant No. 2019YFA0308200).
@@ Appendix A The analytical solution of four-node boson-SGF model with arbitrary gauge phase
The single-excitation dynamics of the four-node boson-SGF model under a symmetric gauge can be analytically solved for arbitrary synthetic flux $\theta_{4}$. Here, we denote the nearest-neighboring hopping phase as $\theta=\theta_{4}/4$, then the four-node boson-SGF Hamiltonian in the single-excitation subspace reads
@@
Here the single-excitation basis kets are $\{|j\rangle\equiv a^{\dagger}_{j}|G\rangle\}$, with $j=1,2,3,4$. The eigenvalues $\lambda_{k}$ and corresponding eigenvectors $\mathbf{f}^{k}$ of $\mathcal{H}^{4}_{\mathrm{SGF}}$ are given by
@@
By setting the initial state as $|\psi(0)\rangle=a_{1}^{\dagger}|G\rangle=(1,0,0,0)^{T}$, the system’s dynamics is determined by Schrödinger equations, with $|\psi(t)\rangle=e^{-iH^{4}_{\rm SGF}t}|\psi(0)\rangle$. We denote that $|\psi(t)\rangle=\sum_{j}C^{j}_{e}(t)|j\rangle$, and the probability amplitudes can be written as
@@
where $f_{j}^{k}=\langle j|\mathbf{f}^{k}\rangle$. Substituting the eigenvalues and eigenvectors from Eq. (24) into the above equation, the population $P^{j}_{e}(t)\equiv|C^{j}_{e}(t)|^{2}$ can be obtained:
@@
The populations of the 2-nd and 4-th nodes are always the same, which reveals the same dynamical process, as shown in Figs. 1(e) and 1(f). In particular, when $\theta_{4}=\pi$ and $\theta=\pi/4$, the 3-rd node can be considered as a dark site whose population vanishes all the time, with $P^{3}_{e}(t,\theta_{4}=\pi)=0$ as shown in Fig. 1(e). Intuitively, the transition amplitude of $1\rightarrow 3$ process can be written as $t_{13}\propto(\exp(i\pi/2)+\exp(-i\pi/2))=0$, with destructive interference from two paths.
@@ Appendix B The dynamics of $n$-node networks
In this part, we study the dynamics of $n$-node networks and show the destruction of chiral flows for the networks whose node number is larger than four, i.e., $n\geq 4$.
We consider the $n$-node boson-SGF model that each node has a $\pi/2$ hopping phase from its next node, and the Hamiltonian reads
@@
Figure 9: Evolution of excited-state population $P_{e}^{j}(t)$ of nodes in (a) five-node and (b) six-node boson-SGF models, respectively. No chiral flow appears regardless of whether the synthetic flux breaks TR symmetry.
@@
with energy spectrum given by $E_{m}=-2J_{0}\sin k_{m}$, where $k_{m}=2\pi m/n$ are the wave numbers within the first Brillouin zone (BZ). For instance, when $n=4$, the wave numbers with $m=0,2,-1$ and $1$ reproduce the four eigenstates in Eq. (24) respectively. We refer to these eigenstates as chiral modes since they have chiral phase shifts $2\pi m/n$ between nearest sites. Note that the system has an internal chiral symmetry $\mathcal{C}$, with $\mathcal{C}^{-1}\mathcal{H}_{\mathrm{SGF}}^{n}\mathcal{C}=-\mathcal{H}_{\mathrm{SGF}}^{n}$, resulting in a symmetric energy spectrum [58]. The chiral operators for even and odd numbers of nodes are given by
@@
As shown in Figs. 9(a) and 9(b), we plot the dynamics of five-node and six-node boson-SGF models and find the perfect chiral flow cannot be supported within these networks even when the hopping phase is $\pi/2$. To understand the underlying reasons, we categorize the boson-SGF networks into odd and even classes based on the value of $n$. When $n$ is odd, the presence of chiral symmetry gives rise to a single zero-energy state, $|\varphi_{0}\rangle=\sum_{j}^{n}a_{j}^{\dagger}|G\rangle/\sqrt{n},~{}E_{0}=0$, corresponding to $k_{m}=0$. Here, we assume that the excitation is initially on the 1-st node. Then using Eq. (25), the evolution of excitation on $j$-th node $P_{e}^{j}(t)=|C_{e}^{j}(t)|^{2}$ can be obtained, with the probability amplitude given by
@@
Figure 10: The energy spectrum of $n$-node boson-SGF model for $n$ ranges from 4 to 11. The dashed lines mean the degenerate energies with two degenerate eigenstates.
Appendix C The analytical solution of four-node and six-node boson-aSGF models
Since the four-node and six-node boson-aSGF models have simple explicit expressions of eigenstates and energy levels, we here derive the analytic solutions for the dynamics of the node excited-state populations $P_{e}^{j}(t)$. In the single-excitation subspace, the Hamiltonian (3) can be written as the following matrix form,

### v2

@@ VI.2 Emitters in magnetic photonic lattices
which contains only the long-range dipole-dipole interaction. The complex hopping amplitude is given by $\tilde{J}_{jk}=g^{2}e^{i\theta_{jk}}G_{0}(\bm{r_{j}},\bm{r}_{k})/(\omega_{e}-\omega_{\ell=0})$, where $G_{0}(\bm{r_{j}},\bm{r}_{k})$ denotes the spatial component of the photonic Green’s function [59]. By setting the gauge-invariant flux $\theta_{4}\equiv\theta_{12}+\theta_{23}+\theta_{34}+\theta_{41}=2\pi$ to break the TR symmetry of each subtriangle plaquette, this effective dipole-dipole interaction can to some extent simulate the four-node spin-aSGF model. Figure  8(b) shows the chiral dynamics of the emitter excitation, where two factors contribute to the imperfection of this chiral flow. First, in this uniform magnetic photonic lattice, the coupling parameter between the auxiliary emitter and node emitter will be $\beta_{c}=1.48$ instead of $2$. Consequently, this deviation prevents the generation of a perfect four-node chiral flow. Second, the unavoidable NNN hopping between emitters from diagonal sites also reduces the fidelity of the chiral flow transmission. Nevertheless, by introducing external coupling channels [18, 28, 13, 82] to eliminate these effects, it is still possible to realize a perfect chiral flow between emitters in this system.
VI.3 Ultracold atoms
The synthetic gauge field can also be found in many experiments with neutral atomic gases. Here, we first describe the general proposal for synthesizing the hopping phase in an atom-light coupling system, where the induced complex couplings between two-level atoms can be used to simulate the spin-SGF model. Then, we introduce the schemes for generating synthetic gauge fields in cold-atom systems, which are quite different from the two-level-system case and allow us to create chiral flows of bosonic atoms. The flexibility of these schemes for generating synthetic gauge fields between atoms also enables the promising implementation of aSGF models in atomic systems.
VI.3.1 External driving fields with two-level atoms
Transitions between atomic levels are usually controlled by external driving fields. One can easily obtain the hopping of a single photon with a certain phase by applying an off-resonant monochromatic field on two separated atoms. For two atoms coupling with a laser field, the full Hamiltonian has the following form:
$H/\hbar=\omega_{c}a^{\dagger}a+\sum_{i=1,2}\omega_{i}\sigma_{i}^{+}\sigma_{i}^{-}+\Omega(\sigma_{i}^{+}ae^{i\phi_{i}}+\mathrm{H.c.}),$ (24)
where $\omega_{i}$ ($i=1,2$) and $\omega_{c}$ are the frequencies of the atoms and the laser field, respectively. The atom-light coupling strength $\Omega$ denotes the Rabi frequency, and $\phi_{i}$ is the coupling phase imparted by the applied field. Spatially separated atoms experience distinct coupling phases due to position-dependent variations in the driving laser field, resulting in a phase difference $\phi=\phi_{1}(\bm{r}_{1})-\phi_{2}(\bm{r}_{2})$. For simplicity, we assume that both atoms share the same transition frequency ($\omega_{1}=\omega_{2}=\omega_{0}$), detuned from the laser frequency by $\Delta=\omega_{0}-\omega_{c}$. In the large-detuning regime where $\Delta\gg\Omega$, the atoms are virtually excited by the driving field. In the interaction picture with $H_{0}=\hbar\omega_{c}(a^{\dagger}a+\sigma_{1}^{+}\sigma_{1}^{-}+\sigma_{2}^{+}\sigma_{2}^{-})$ and $V=H-H_{0}$, the Hamiltonian becomes
$\widetilde{H}/\hbar=$ $e^{iH_{0}t}Ve^{-iH_{0}t}$ (25)
$=$ $\sum_{i=1,2}\Delta\sigma_{i}^{+}\sigma_{i}^{-}+\Omega(\sigma_{i}^{+}ae^{i\phi_{i}}+\mathrm{H.c.}).$
To obtain the effective Hamiltonian of the dynamics, we perform a Schieffer-Wolff transformation such that $\widetilde{H}_{\mathrm{eff}}=e^{S}\widetilde{H}e^{-S}$, with $S=\sum_{i}\frac{\Omega}{\Delta}(\sigma_{i}^{-}a^{\dagger}-\sigma_{i}^{+}a)$. We can then eliminate the degree of freedom of the light, and the atoms are thus decoupled from the light field. The effective Hamiltonian is given by
$\widetilde{H}_{\mathrm{eff}}/\hbar=\sum_{i=1,2}(\Delta+\lambda)\sigma_{i}^{+}\sigma_{i}^{-}+\lambda e^{i\phi}(\sigma_{1}^{+}\sigma_{2}^{-}+\mathrm{H.c.}),$ (26)
where $\lambda=\Omega^{2}/\Delta$ denotes the coupling strength and Lamb shift of atoms due to a second-order (virtual photon-exchange) process from the off-resonant atom-light coupling. The effective Hamiltonian indicates a complex-valued atom-atom interaction from the dynamics. The phase attached to the field now appears in the photon-exchange interaction between atoms. In this sense, giving different dressing fields to any two neighboring atoms induces different phases of excitation hopping. For $N$ atoms with monochromatic dressing, the system’s Hamiltonian is given by $H=\sum_{k=1}^{N-1}ge^{i\theta_{k}}\sigma_{k}^{+}\sigma_{k+1}^{-}+\mathrm{H.c.}$ If atoms form a closed coupling loop, the total phase—i.e., the synthetic flux—becomes a nontrivial quantity, which is gauge invariant under any gauge transformations. Since the two-level atom is described by Pauli spin operators, this process can be exploited to create chiral flows of spin.
VI.3.2 Noninertial frame of reference and geometric gauge potentials with bosonic atoms
Constructing the boson-SGF model entails the hopping of bosonic atoms instead of excitation. For charge-neutral atoms, SGFs are obtained in a noninertial frame of reference, as in rotating systems [83, 84, 85, 86], or by inducing geometric phases [87, 88, 89, 90, 91]. In these scenarios, the center-of-mass motion of the atoms acquires Peierls phases that are linked to the spatial gauge fields within the system. Consider an atom of mass $M$ trapped in a potential field $V(\bm{r})$, where $\bm{r}$ denotes the center-of-mass position vector. The atomic Hamiltonian is then expressed as
$H_{\mathrm{at}}=\frac{\bm{p}^{2}}{2M}+V(\bm{r}).$ (27)
If the system rotates about the direction of a unit vector $\hat{\mathbf{n}}$ at a constant angular frequency $\bm{\mathcal{W}}=\mathcal{W}\hat{\mathbf{n}}$, the position vector in the spatially rotating frame becomes $\bm{r}\rightarrow\bm{r}^{\prime}=\mathcal{D}_{\mathbf{\hat{n}}}(t)\bm{r}\mathcal{D}_{\mathbf{\hat{n}}}^{\dagger}(t)$, where $\mathcal{D}_{\mathbf{\hat{n}}}(t)=\mathrm{exp}{(-i\bm{\mathcal{W}}\cdot\bm{L}t/\hbar)}$ is the rotation operator and $\bm{L}=\bm{r}\times\bm{p}$ is the angular momentum operator. To obtain a time-independent Hamiltonian, we work in the rotating frame such that [8]
$H=$ $\mathcal{D}_{\mathbf{\hat{n}}}^{\dagger}H_{\mathrm{at}}(\bm{r^{\prime}}(t))\mathcal{D}_{\mathbf{\hat{n}}}-i\hbar\mathcal{D}_{\mathbf{\hat{n}}}^{\dagger}\frac{\partial}{\partial t}\mathcal{D}_{\hat{\mathbf{n}}}$ (28)
$=$ $\frac{(\bm{p}-\bm{A})^{2}}{2M}+V(\bm{r})+W_{\mathrm{cen}}(\bm{r}),$
[…]
In addition to engineering synthetic gauge fields in noninertial reference frames, employing geometric gauge potentials offers an alternative approach. Crucially, the internal degrees of freedom of atomic states provide a tool for creating geometric phases. Considering the internal states of atoms, the full Hamiltonian is now written as
$H_{\mathrm{at}}=\left[\frac{\bm{p}^{2}}{2M}+V(\bm{r})\right]\mathbb{I}+\mathbb{M}(\bm{r}),$ (29)
where $\mathbb{I}$ is the identity operator in the internal Hilbert space. The coupling term $\mathbb{M}(\bm{r})$ between atomic internal states and the external field characterizes the dynamics of the atom’s internal degrees of freedom, and its spatial dependence links internal states with the atomic position. The atomic state vector can be expanded in the dressed-state basis $|\chi_{m}\rangle$ as
$|\psi(\bm{r},t)\rangle=\sum_{m=1}^{N}\phi_{m}(\bm{r},t)\left|\chi_{m}(\bm{r})\right\rangle.$ (30)
In the basis of dressed states, the coupling term $\mathbb{M}(\bm{r})$ becomes a diagonal matrix $\hat{\varepsilon}$ with eigenenergies $\varepsilon_{m}$, and the Hamiltonian in this representation reads [87, 8]
$\hat{H}=\frac{(\bm{p}-\bm{A})^{2}}{2M}+V(\bm{r})+\hat{\varepsilon}.$ (31)
This implies that atomic external motion is affected by a geometric gauge field, whose matrix elements are given by $\langle\chi_{n}|\bm{A}|\chi_{m}\rangle=i\hbar\langle\chi_{n}|\nabla\chi_{m}\rangle$. The geometric potential arises from the interplay between atomic external (center-of-mass) dynamics and its internal degree of freedom. Since the dressed states vary with atomic position, the momentum operator that acts on these position-dependent states effectively couples the internal Hilbert space. Suppose that nonadiabatic transitions between dressed states are significantly suppressed due to the large energy gap; in this case, we can project the system’s dynamics onto a certain atomic ground-state manifold. If an atom is initially prepared in the internal state $|\chi_{1}\rangle$, by substituting Eq. (30) into the time-dependent Schrödinger equation, we obtain the equation of external motion [7],
$i\hbar\frac{\partial\phi_{1}(\bm{r})}{\partial t}=\left[\frac{(\bm{p}-\bm{A})^{2}}{2M}+V(\bm{r})+\varepsilon_{1}+W\right]\phi_{1}(\bm{r}).$ (32)
Now the reduced vector potential is denoted by $\bm{A}=i\hbar\langle\chi_{1}|\nabla\chi_{1}\rangle$. This is associated with the Born-Oppenheimer approximation, where the external dynamics are separated from the internal state. Here, the additional scalar potential $W(\bm{r})={\hbar^{2}}\sum_{m=2}^{N}|\langle\chi_{m}|\chi_{1}\rangle|^{2}/{2M}$ appears due to the nonadiabatic transitions from other dressed states, which can be absorbed in the potential field $V(\bm{r})$.
In experiments, the $\Lambda$-type atomic-level structure is one of the most widely used configurations for generating light-induced gauge fields, in which the internal states of atoms adiabatically follow the dark state, as mentioned above. The dark state is formed by a linear combination of two ground states, and it suppresses the heating problem caused by spontaneous emission from the excited state. As shown in Fig. 9(a), the light-atom coupling term is now given by
Figure 9: (a) Level structure of $\Lambda$-type atom with a dark state determined by Rabi frequencies. (b) Raman dressing scheme for Zeeman sublevels of the ${}^{87}\mathrm{Rb}$ atom.
$\mathbb{M}=\hbar\delta(|g_{2}\rangle\langle g_{2}|-|g_{1}\rangle\langle g_{1}|)+\sum_{i=1}^{2}\hbar(\Omega_{i}|g_{i}\rangle\langle e|+\mathrm{H.c.}),$ (33)
where the two laser beams induce coherent transitions between ground states $|g_{i}\rangle$ and the excited state $|e\rangle$, with position-dependent Rabi frequencies $\Omega_{i}(\bm{r})=|\Omega_{i}|e^{i\phi_{i}}$ ($i=1,2$). When the excited state $|e\rangle$ is in resonance, i.e., $\delta=0$, the system manifests a dark state $|D\rangle=(\Omega_{2}|g_{1}\rangle-\Omega_{1}|g_{2}\rangle)/\Omega$, and $\Omega=\sqrt{|\Omega_{1}|^{2}+|\Omega_{2}|^{2}}$. For atoms initially prepared in this dark state, the associated geometric gauge potential is given by [7, 8]
$\begin{gathered}\bm{A}=i\hbar\langle D(\bm{r})|\nabla D(\bm{r})\rangle=-\hbar\frac{|\zeta|^{2}\nabla\phi}{1+|\zeta|^{2}},\\ \zeta=\Omega_{1}(\bm{r})/\Omega_{2}(\bm{r})=|\zeta|\mathrm{exp}(i\phi),\end{gathered}$ (34)
where $|\zeta|=|\Omega_{1}|/|\Omega_{2}|$ and $\phi=\phi_{1}-\phi_{2}$ denote the amplitude ratio of two Rabi frequencies and the phase difference, respectively. In addition, spatial gradients of these two parameters must be nonzero to induce a nontrivial synthetic flux along a closed loop of atomic motion. This can be realized using two laser beams with orbital angular momentum [97]. For bosonic ultracold atoms hopping between individual traps, the light-induced SGF can be engineered to realize a boson-SGF model with chiral flows. Alternatively, one can apply the Raman dressing scheme of a ladder-type atom where the excited state is replaced with another ground state, as sketched in Fig. 9(b). This configuration can be realized with alkali ultracold atoms, like the Raman-dressed ${}^{87}\mathrm{Rb}$ Bose-Einstein condense in the $F=1$ ground-state manifold [98]. To create Zeeman sublevels $|m_{F}=0,\pm 1\rangle$, the degeneracy is lifted by an external bias magnetic field. The spatially varying two-photon detuning $\delta$ in the Raman coupling generates a synthetic vector potential for the atoms, leading to position-dependent eigenstates. This Raman dressing scheme can also be applied to a two-level system; however, intrinsic limitations arise concerning the spatial range of SGFs and the balance between off-resonant scattering and Raman coupling [91].
@@ VII Conclusion and Outlook
In conclusion, we investigate the boson-SGF model and show the chiral excitation flow in the three-node case with a broken TR symmetry; however, the pure SGF model cannot drive the chiral dynamics when it extends to the four-node case. By introducing the auxiliary node to couple to network nodes, we recreate the chiral flow in the four-node boson-aSGF network. The excitation dynamics of this four-node boson-aSGF network suggest that it can exhibit chiral flow only when the synthetic flux of each triangular substructure breaks TR symmetry, i.e., $\theta_{\vartriangle}\neq p\pi$ ($p\in\mathbb{Z}$). This links the well-studied three-node chiral flow and the boson-aSGF model for creating perfect chiral flows in multinode networks, where these triangular substructures are also retained, with $(p+1/2)\pi$ ($p\in\mathbb{Z}$) synthetic flux. When we generalize this framework to the spin or two-level atomic system where TR symmetry is preserved, it demonstrates opposite chiral dynamics for the spin-up and spin-down states. Further, we introduce a universal scheme for constructing $n$-node chiral-flow Hamiltonians with the aSGF model and provide illustrative examples for odd and even $n$ values. During this process, we find the criteria for a Hamiltonian to support the chiral flow in a specific excitation subspace, which indicates the requirement for a symmetric and equally spaced energy spectrum and a complete set of chiral modes. Our framework for finding chiral flows applies to various systems such as bosons, spins, and two-level atoms. It is also applicable to different types of interactions, including the NN complex hopping forming a two-body Hamiltonian, as well as the three-body interactions like SCI and ASI in spin systems. Nonetheless, to realize the chiral flow in a larger network, the number of required hoppings scales as $n^{2}/2$, showing significant complexity growth. Therefore, we propose a ladder network with the complexity scaling as $n$. This can serve as a more practical scheme for unidirectional transmission between remote emitters. In addition, the imperfections of networks are also discussed, where the fidelity of chiral flow shows robustness in the presence of hopping disorder but is sensitive to the inhomogeneity of node (boson mode or atomic) frequency.
@@
The multinode aSGF model and the extension of the four-node network can be realized with superconducting circuits by employing Floquet modulation, offering a practical approach to implementing perfect state transfer in circuit QED systems. The magnetic photonic lattice can also mimic this model to a certain extent with atomic emitters. Moreover, synthetic gauge fields in cold-atom systems provide a feasible route for implementing both boson- and spin-aSGF models. The experimental feasibility of our scheme, along with its tolerance to imperfections, makes it a practical and promising approach for realizing chiral state transfer and unidirectional entanglement distribution. We also expect that our proposal could be used to develop nonreciprocal quantum devices [75], such as $n$-node circulators, and benefit future quantum technologies.
@@ VIII Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 12375025 and No. 11874432) and the National Key R&D Program of China (Grant No. 2019YFA0308200).
@@ Appendix A The analytical solution of four-node boson-SGF model with arbitrary gauge phase
The single-excitation dynamics of the four-node boson-SGF model under a symmetric gauge can be analytically solved for arbitrary synthetic flux $\theta_{4}$. Here, we denote the nearest-neighboring hopping phase as $\theta=\theta_{4}/4$. The four-node boson-SGF Hamiltonian in the single-excitation subspace then reads
@@
Here, the single-excitation basis kets are $\{|j\rangle\equiv a^{\dagger}_{j}|G\rangle\}$, with $j=1,2,3,4$. The eigenvalues $\lambda_{k}$ and corresponding eigenvectors $\mathbf{f}^{k}$ of $\mathcal{H}^{4}_{\mathrm{SGF}}$ are given by
@@
By setting the initial state as $|\psi(0)\rangle=a_{1}^{\dagger}|G\rangle=(1,0,0,0)^{T}$, the system’s dynamics are determined by the Schrödinger equation, with $|\psi(t)\rangle=e^{-iH^{4}_{\rm SGF}t}|\psi(0)\rangle$. We denote $|\psi(t)\rangle=\sum_{j}C^{j}_{e}(t)|j\rangle$, and the probability amplitudes can be written as
@@
where $f_{j}^{k}=\langle j|\mathbf{f}^{k}\rangle$. Substituting the eigenvalues and eigenvectors from Eq. (36) into the above equation, the population $P^{j}_{e}(t)\equiv|C^{j}_{e}(t)|^{2}$ can be obtained [64]:
@@
The populations of the second and fourth nodes are always the same, which reveals the same dynamical process, as shown in Figs. 1(e) and 1(f). In particular, when $\theta_{4}=\pi$ and $\theta=\pi/4$, the third node can be considered as a dark site whose population vanishes all the time, with $P^{3}_{e}(t,\theta_{4}=\pi)=0$, as shown in Fig. 1(e). Intuitively, the transition amplitude of the $1\rightarrow 3$ process can be written as $t_{13}\propto(\exp(i\pi/2)+\exp(-i\pi/2))=0$, with destructive interference from two paths.
@@ Appendix B The dynamics of $n$-node networks
In this appendix, we study the dynamics of $n$-node networks and show the destruction of chiral flows for the networks whose number of nodes is greater than four, i.e., $n\geq 4$.
We consider the $n$-node boson-SGF model in which each node has a $\pi/2$ hopping phase from its next node, and the Hamiltonian reads
@@
Figure 10: Evolution of excited-state population $P_{e}^{j}(t)$ of nodes in (a) five-node and (b) six-node boson-SGF models. No chiral flow appears, regardless of whether the synthetic flux breaks TR symmetry.
@@
with an energy spectrum given by $E_{m}=-2J_{0}\sin k_{m}$, where $k_{m}=2\pi m/n$ are the wave numbers within the first Brillouin zone (BZ). For instance, when $n=4$, the wave numbers with $m=0,2,-1$ and $1$ reproduce the four eigenstates in Eq. (36). We refer to these eigenstates as chiral modes since they have chiral phase shifts $2\pi m/n$ between nearest sites. Note that the system has an internal chiral symmetry $\mathcal{C}$, with $\mathcal{C}^{-1}\mathcal{H}_{\mathrm{SGF}}^{n}\mathcal{C}=-\mathcal{H}_{\mathrm{SGF}}^{n}$, resulting in a symmetric energy spectrum [67]. The chiral operators for even and odd numbers of nodes are given by
@@
In Figs. 10(a) and 10(b), we plot the dynamics of five-node and six-node boson-SGF models, finding that the perfect chiral flow cannot be supported within these networks, even when the hopping phase is $\pi/2$. To understand the underlying reasons for this, we categorize the boson-SGF networks into odd and even classes based on the value of $n$. When $n$ is odd, the presence of chiral symmetry gives rise to a single zero-energy state, $|\varphi_{0}\rangle=\sum_{j}^{n}a_{j}^{\dagger}|G\rangle/\sqrt{n},~{}E_{0}=0$, corresponding to $k_{m}=0$. Here, we assume that the excitation is initially on the first node. Then, using Eq. (37), the evolution of excitation on $j$th node $P_{e}^{j}(t)=|C_{e}^{j}(t)|^{2}$ can be obtained, with the probability amplitude given by
@@
Figure 11: Energy spectrum of $n$-node boson-SGF model for $n$ values from 4 to 11. The dashed lines indicate degenerate energies with two degenerate eigenstates.
Appendix C Analytical solutions of four-node and six-node boson-aSGF models
Since the four-node and six-node boson-aSGF models have simple explicit expressions of eigenstates and energy levels, here, we derive the analytic solutions for the dynamics of the node excited-state populations $P_{e}^{j}(t)$. In the single-excitation subspace, the Hamiltonian (3) can be written in the following matrix form with $n=4$: